OpenStudy (anonymous):

You buy a new car for $32,000. The value of the car decreases by 12% each year. a. Write a rule for the average yearly value of the car (in dollars) in terms of the current year. Let n=the current year. b. In about how many years will the value of the car fall to $13,077? I think this was a geometric sequence and so I used : an = a1(r)^(n-1) an = 32000(-.12)^(n-1) and this is my rule. Then I put 13077 in an and solved 13077 = 32000(-.12)^(n-1) 13077/32000 = -.12^(n-1) log 13077/32000 = log-.12^(n-1) log13077/32000/log-.12 = n – 1 log13077/32000/log-.12 +

OpenStudy (anonymous):

@dpaInc pl help..the quest been up for 5 hrs nw..

OpenStudy (anonymous):

sorry man... i really can't think anymore... it's 1 am here... i'm gonna get some shut eye...

OpenStudy (anonymous):

its ok..thankx for comming tho.

OpenStudy (anonymous):

i'll check up on this tomorrow...

OpenStudy (anonymous):

tankx again ;) @kropot72 any hpe?

OpenStudy (kropot72):

Sorry, I need time to think on this one.

OpenStudy (anonymous):

itz okay tak ur tym

OpenStudy (tommo_lcfc):

I can help you! Now the first thing to note is that if the price decreases by 12% each year, this is equivalent to 88% each year (100-12 = 88). As it decreases, the ratio is 0.88. If it was increasing by 12%, the ratio would be 1.12. Common ratios can never be negative. So the nth term is 32000 x 0.88^(n-1). Equate this to 13077. So 13077 = 32000 x 0.88^(n-1) 13077/32000 = 0.88^(n-1) log(13077/32000) = (n-1) log (0.88) n-1 = log(13077/32000) / log(0.88) = 7.000370328. So n = 8.000370328 = 8 years. Do you get that?